Keeping rolling a 4-sided die until you see the first ACE(SPOT 1). Let X be the number of rolls resulting

1. Find P[X = 3]

2. Find P[X >= 3]

3. Find E[X]

4. Find V[X]
Answer #1

here this follows geometric distribution with parameter p=probability of ace =1/4 =0.25

1)

probability = P(X=3)= (1-0.25)3-1*0.25= 0.1406

2)

probability = P(X>=3)= 1-P(X<=2)= (1-0.25)3-1= 0.5625

3)

here mean of distribution=E(x)=1/p= 4.00

4)

varaince =σ2=(1-p)/p2= 12.00